40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 2 02:44:45 2014, Jenne, Update, LSC, Tried cav pole compensation trick - fail 
    Reply  Wed Dec 3 01:46:12 2014, Koji, Update, LSC, Tried cav pole compensation trick - fail 
Message ID: 10746     Entry time: Tue Dec 2 02:44:45 2014     Reply to this: 10748
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Tried cav pole compensation trick - fail 

[Jenne, Diego]

First, random notes:

  • saw a violin peak in CARM / DARM at 638.0Hz.  Assumed it was one of the ETMs, even though it doesn't match any of the frequencies in our handy-dandy chart: wiki resonances
    • Put an extra notch in the ETM violin filters.
    • Just now realized that I was actuating MC2 at the time for CARM (although 638 is also not what we have in the chart for MC2).  The MC2/ETM violin filters should be shared between eachother.
  • Measured CARM and DARM loops on ALS comm and diff, gains should be 8, not 6.  Fixed in Lock_ALS_CARM_and_DARM script. 
  • MC has been fussy tonight.  I started actuating CARM on ETMs, and that helped, but we've still had several unexplained MC locklosses. 
    • PC and FSS Slow are okay right now, but they have been mad earlier tonight.  Do we need to check the PID tuning for FSS slow?
  • When I first started locking this evening, I was able to hold nice high arm powers (with the usual factor of 2+ RIN), so the IFO seemed okay except for the fussy MC.

Koji suggested last week that we put a cavity pole filter into the ALS error signals, and then compensate for that in the CARM and DARM servos.  The idea is that any RF signals we want to transfer to will have some kind of frequency dependence, and at the final zero CARM offset that will be a simple cavity pole. 

I put a pole at 200 Hz, with a zero at 6 kHz into the LSC-ALS[X,Y] filter banks in FM1, and then also put a zero at 200 Hz with a pole at 6 kHz into both the CARM and DARM servos at FM7.  Ideally I wouldn't have the 6kHz in there, but the compensation filter in the CARM/DARM servos needs a pole somewhere, so I put in the zero in the ALS signals so that they match.  Foton thinks that multiplying the two filters should give a flat response, to within 1e-6dB and 1e-6 deg. 

We can lock CARM and DARM on ALS with the new filters, but it seems to be not very stable.  We've measured transfer functions in both configurations, and between 50-500Hz, there is no difference (i.e., our matching filters are matching, and cancelling each other out).  We sometimes spontaneously lose lock when we're just sitting somewhere with the new configuration, and we cannot run any find IR resonance scripts and stay locked.  We've tried the regular old script, as well as Diego's new gentler script.  We always fail with the regular script during the coarse scan.  With Diego's script, we made it through the coarse scan, but spontaneously lost lock while the script was calculating the location of the peak.  So, we determine that there is something unstable about the new configuration that we don't understand.  Turning off all the new filters and going back to the old configuration is just as robust as always.  Confusing. 

 

ELOG V3.1.3-